# Robust Implicit Regularization via Weight Normalization

### v3

@@ 3 Proofs
Here is the outline of the proofs. First we prove the basic properties of the dynamics (13) and (14), including scaling (Lemma 3.1, constant norm (Lemma 3.2), constant sign (Lemma 3.3), and decreasing loss (Lemma 3.5).
We then prove Theorem 2.2 via the comparison (Lemma 3.9) between invariants without weight normalization (Lemma 3.7) and invariants with weight normalization (Lemma 3.8), together with the boundedness (Lemma 3.10) and the uniqueness (Lemma 3.11) of the trajectories.
We further prove Theorem 2.3 based on the additional boundedness (Lemma 3.12) for $L=1,2$. We prove Theorem 2.4 through a different approach than the previous parts.
Finally, we analyze the trajectories through dynamic reduction (Lemma 3.14) and the non-increasing Bregman divergence (Lemma 3.18), which is similar to the proof in [8].
@@ 3.1 Basic properties
$\nabla\mathcal{L}({\bf x})$ $=[{\bf A}^{\top}({\bf A}{\bf x}^{\odot L}-{\bf b})]\odot({\bf x}^{\odot L-1}),$ (22)
$\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u})$ $=\frac{{\bf u}^{\top}}{\|{\bf u}\|_{2}}\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right),$ (23)
$\nabla_{{\bf u}}\widetilde{\mathcal{L}}(r,{\bf u})$ $=\frac{r}{\|{\bf u}\|_{2}}(I-\mathcal{P}_{{\bf u}})\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right).$ (24)
@@
Suppose $(r,{\bf u})$ follows (13) and (14) with initialization $(r_{0},{\bf u}_{0})$ and learning rate $(\eta_{r},\eta_{\bf u})$. Fix ${a}>0$. Suppose $(r^{({a})},{\bf u}^{({a})})$ follows (13) and (14) with initialization $(r_{0},{a}{\bf u}_{0})$ and learning rate $(\eta_{r},{a}^{2}\eta_{\bf u})$. Then $r^{({a})}(t)=r(t)$ and ${\bf u}^{({a})}(t)={a}{\bf u}(t)$ for all $t\geq 0$. As a result,
@@
First note that for
$r^{({a})}=r\quad\text{and}\quad{\bf u}^{({a})}={a}{\bf u}$ (25)
we obtain
Consequentially,
$\partial_{t}r^{({a})}$ $=-\eta_{r}\nabla_{r^{({a})}}\widetilde{\mathcal{L}}(r^{({a})},{\bf u}^{({a})})=\partial_{t}r,$ (26)
$\partial_{t}{\bf u}^{({a})}$ $=-{a}^{2}\eta_{\bf u}\nabla_{{\bf u}^{({a})}}\widetilde{\mathcal{L}}(r^{({a})},{\bf u}^{({a})})={a}\partial_{t}{\bf u}.$ (27)
Since (25) holds for $t=0$ and is preserved due to (26) and (27), (25) holds for all $t\geq 0$. ∎
@@
A direct computation yields
because ${\bf u}^{\top}(I-\mathcal{P}_{{\bf u}})=0$. This implies the claim. ∎
@@
Note that $\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u})$ is local Lipschitz continuous in $r$, and $\nabla_{{\bf u}}\widetilde{\mathcal{L}}(r,{\bf u})$ is local Lipschitz continuous in ${\bf u}$. Hence by the Picard–Lindelöf theorem the trajectory is unique. In particular, $\partial_{t}r$ and $u_{n}$ cannot reach zero at finite time, since this would otherwise contradict to the uniqueness of the trajectory (if we apply Picard–Lindelöf theorem backward in time). We now show that $r=0$ implies that $\partial_{t}r=0$ and $u_{n}=0$ implies that $\partial_{t}u_{n}=0$. By (22), (23), and (24), we have
$\partial_{t}r$ $=-\eta_{r}\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u})=-\frac{\eta_{r}}{\|{\bf u}\|_{2}}{\bf u}^{\top}\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right)$ (28)
$=-\eta_{\bf u}\frac{r}{\|{\bf u}\|_{2}}\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right)-\eta_{\bf u}\frac{r\partial_{t}r}{\eta_{r}\|{\bf u}\|_{2}^{2}}{\bf u}.$ (29)
@@
Note that although Lemma 3.3 shows that the ${\bf u}(t)>0$ for all $t\geq 0$, it does not guarantee that its limit is strictly positive. Hence later on while discussing the limit we will need some slightly stronger results (Lemma 3.10 and the Lemma 3.11).
We now examine the dynamics of the loss, which is non-increasing in time.
@@
If $(r,{\bf u})$ follows (13) and (14), then the loss ~$\widetilde{\mathcal{L}}(r,{\bf u})$ is non-increasing in time, i.e. $\partial_{t}\widetilde{\mathcal{L}}(r,{\bf u})\leq 0$.
@@
By chain rule we have
This completes the proof. ∎
We can even derive the convergence rate under the assumption that some entries are uniformly lower bounded, which in fact holds if the trajectory is upper bounded, as we will later on prove in Lemma 3.10 in Section 3.2.
@@
Let ${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$ and $(r,{\bf u})$ follows (13) and (14) with $\|{\bf u}_{0}\|_{2}^{2}=1$. Let $t_{0}\geq 0$. Suppose there exists constant $c_{r},c_{\bf u},c_{\bf x}>0$ such that $\eta_{r}\geq c_{r}$, $\eta_{\bf u}\geq c_{\bf u}$, and $|{\bf x}_{I}|\geq c_{\bf x}$ for all $t\in[t_{0},T]$, for some index set $I\subset[N]$. Denote ${\bf A}_{I}$ to be the sub-matrix of ${\bf A}$ with columns indexed by $I$. If ${\bf A}_{I}$ has full rank, then for all $t\geq t_{0}$,
$\mathcal{L}({\bf x}(t))\leq\mathcal{L}({\bf x}(t_{0}))\exp\left(-\min\left(c_{r},c_{\bf u}c_{\bf x}^{2}|I|\right)2Lc_{\bf x}^{2}\sigma_{\min}^{2}({\bf A}|_{I})(t-t_{0})\right)$ (30)
@@
By Lemma 3.2, $\|{\bf u}\|_{2}=\|{\bf u}_{0}\|_{2}=1$. By Lemma 3.5, we have
Note that by (23) and (24),
On the other hand,
Putting together the estimates, we obtain
By Gronwall’s inequality we get the linear convergence rate
This completes the proof. ∎
@@ 3.2 Invariants of the flow
Let $L\in\mathbb{N}$. Suppose ${\bf x}$ follows the dynamics in (11). Then the quantity
$h_{0}(t):=\left(I-{\bf A}^{\dagger}{\bf A}\right)\cdot\begin{cases}\log({\bf x}(t))&\text{if }L=2\\ {\bf x}(t)^{\odot 2-L}&\text{if }L\neq 2\end{cases}$ (31)
@@
Let $L\in\mathbb{N}$, ${\tilde{\eta}}>0$, and $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$. Suppose $r,{\bf u}$ follow the dynamics in (13) and (14) with $\|{\bf u}_{0}\|_{2}=1$. Then the quantity
$h_{\tilde{\eta}}(t):=(I-{\bf A}^{\dagger}{\bf A})\cdot\begin{cases}\log\left({\bf u}(t)\exp\left(\frac{1}{2{\tilde{\eta}}}r(t)^{2}\right)\right)&\text{if }L=2\\ {\bf u}(t)^{\odot 2-L}\exp\left(\frac{2-L}{2{\tilde{\eta}}}r(t)^{2}\right)&\text{if }L\neq 2\end{cases}$ (32)
@@
Proof.
Since $\|{\bf u}_{0}\|_{2}=1$, by Lemma 3.2 $\|{\bf u}(t)\|_{2}=1$ for all $t\geq 0$. By substituting $\eta_{r}={\tilde{\eta}}$, $\eta_{\bf u}=1$, and $\|{\bf u}\|_{2}=1$ into the dynamics (29), we obtain
$\partial_{t}{\bf u}=-r\nabla\mathcal{L}\left(r{\bf u}\right)-\frac{r}{{\tilde{\eta}}}{\bf u}\partial_{t}r=-r\nabla\mathcal{L}\left(r{\bf u}\right)-\frac{1}{2{\tilde{\eta}}}\partial_{t}(r^{2}){\bf u}.$ (33)
@@
$(I-\mathcal{P}_{{\bf A}})\left({\bf u}^{\odot 1-L}\odot\nabla\mathcal{L}\left(r{\bf u}\right)\right)=(I-\mathcal{P}_{{\bf A}})\left({\bf A}^{\top}\left({\bf A}\left(r{\bf u}\right)^{\odot L}-{\bf b}\right)r^{L-1}\right)=0,$ (34)
by applying the operation $(I-\mathcal{P}_{{\bf A}})[{\bf u}^{\odot 1-L}\odot\,\cdot\,]$ to (33) we have
$(I-\mathcal{P}_{{\bf A}})({\bf u}^{\odot 1-L}\odot\partial_{t}{\bf u})=-\frac{1}{2{\tilde{\eta}}}\partial_{t}(r^{2})(I-\mathcal{P}_{{\bf A}}){\bf u}^{\odot 2-L}.$ (35)
@@
For $L=2$, since ${\bf u}^{\odot-1}\odot\partial_{t}{\bf u}=\partial_{t}\log({\bf u})$, (35) can be expressed as
@@
$(I-\mathcal{P}_{{\bf A}})\left(\log({\bf u})+\frac{r^{2}}{2{\tilde{\eta}}}\bm{1}\right)=(I-\mathcal{P}_{{\bf A}})\left(\log({\bf u}_{0})+\frac{r_{0}^{2}}{2{\tilde{\eta}}}\bm{1}\right).$ (36)
@@
For $L\neq 2$ the left hand side of (35) can be written as $\frac{1}{2-L}\partial_{t}(I-\mathcal{P}_{{\bf A}}){\bf u}^{\odot 2-L}$. Let $\widetilde{{\bf u}}=(I-\mathcal{P}_{{\bf A}}){\bf u}^{\odot 2-L}$. Then we have that
@@
$\widetilde{{\bf u}}\cdot\exp\left(\frac{2-L}{2{\tilde{\eta}}}r^{2}\right)=\widetilde{{\bf u}}_{0}\cdot\exp\left(\frac{2-L}{2{\tilde{\eta}}}r_{0}^{2}\right)$ (37)
@@
[…]
$\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)\log({\bf x}(t))$ $=\left(I-\mathcal{P}_{{\bf A}}\right)\log({\bf x}(0))$ (38)
[…]
@@
$\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right){\bf x}^{\odot 2-L}(t)$ $=\left(I-\mathcal{P}_{{\bf A}}\right){\bf x}^{\odot 2-L}(0)$ (40)
[…]
@@
$\gamma(r_{0},r)$ $:=\frac{r}{r_{0}}\exp\left(\frac{r_{0}^{2}-r^{2}}{2{\tilde{\eta}}}\right).$ (42)
@@
[…]

### v4

@@ 3 Proofs
In this section we will present the main lemma and theorems, along with some of the key proof techniques. The rest of the proofs can be found in Appendix A. Here is the outline.
  1. 1.
    Basic properties of the dynamics: we use Lemma 3.1 to reduce the number of parameters, Lemma 3.2 to avoid division by zero, Lemma 3.3 to control the signs, Lemma 3.5 and Lemma 3.6 to guarantee that the loss is indeed decreasing.
  2. 2.
    Proof of Theorem 2.2: the key proof technique is to compare the invariants with and without weight normalization (Lemma 3.7, 3.8, 3.9). To ensure the well-posedness of such comparison, we prove the boundedness (Lemma 3.10) and the uniqueness (Lemma 3.11) of the trajectories.
  3. 3.
    Proof of Theorem 2.3: it relies on additional boundedness (Lemma 3.12) for $L=1,2$.
  4. 4.
    Proof of Theorem 2.4: we use techniques similar to the ones in [8], where we analyze the trajectories through dynamic reduction (Lemma 3.14) and the Bregman divergence (Lemma 3.18.
@@ 3.1 Basic properties
$\nabla\mathcal{L}({\bf x})$ $=[{\bf A}^{\top}({\bf A}{\bf x}^{\odot L}-{\bf b})]\odot({\bf x}^{\odot L-1}),$ (3.1)
$\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u})$ $=\frac{{\bf u}^{\top}}{\|{\bf u}\|_{2}}\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right),$ (3.2)
$\nabla_{{\bf u}}\widetilde{\mathcal{L}}(r,{\bf u})$ $=\frac{r}{\|{\bf u}\|_{2}}(I-\mathcal{P}_{{\bf u}})\nabla\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right).$ (3.3)
@@
Suppose $(r,{\bf u})$ follows (1.13) and (1.14) with initialization $(r_{0},{\bf u}_{0})$ and learning rate $(\eta_{r},\eta_{\bf u})$. Fix ${a}>0$. Suppose $(r^{({a})},{\bf u}^{({a})})$ follows (1.13) and (1.14) with initialization $(r_{0},{a}{\bf u}_{0})$ and learning rate $(\eta_{r},{a}^{2}\eta_{\bf u})$. Then $r^{({a})}(t)=r(t)$ and ${\bf u}^{({a})}(t)={a}{\bf u}(t)$ for all $t\geq 0$. As a result,
@@
See Appendix A. ∎
@@
See Appendix A. ∎
@@
See Appendix A. ∎
@@
Note that although Lemma 3.3 shows that the ${\bf u}(t)>0$ for all $t\geq 0$, it does not guarantee that its limit is strictly positive, because hitting zeros will have some undesirable consequence. Hence later on while discussing the limit we will need some slightly stronger results (Lemma 3.10 and the Lemma 3.11). Note that although we will prove that the trajectories is bounded away from zero by some positive constant, this constant can be small when the initialization is small. Hence such bound does not prevent us from getting close to a sparse solution. The need of such constant is rather a technical requirement for the proof.
We can show that the loss is non-increasing. Moreover, we can derive the convergence rate under the assumption that some entries are uniformly lower bounded, which in fact holds if the trajectory is upper bounded, as we will later on prove in Lemma 3.10 in Section 3.2.
@@
If $(r,{\bf u})$ follows (1.13) and (1.14), then the loss ~$\widetilde{\mathcal{L}}(r,{\bf u})$ is non-increasing in time, i.e. $\partial_{t}\widetilde{\mathcal{L}}(r,{\bf u})\leq 0$.
@@
See Appendix A. ∎
@@
Let ${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$ and $(r,{\bf u})$ follows (1.13) and (1.14) with $\|{\bf u}_{0}\|_{2}^{2}=1$. Let $t_{0}\geq 0$. Suppose there exists constant $c_{r},c_{\bf u},c_{\bf x}>0$ such that $\eta_{r}\geq c_{r}$, $\eta_{\bf u}\geq c_{\bf u}$, and $|{\bf x}_{I}|\geq c_{\bf x}$ for all $t\in[t_{0},T]$, for some index set $I\subset[N]$. Denote ${\bf A}_{I}$ to be the sub-matrix of ${\bf A}$ with columns indexed by $I$. If ${\bf A}_{I}$ has full rank, then for all $t\geq t_{0}$,
$\mathcal{L}({\bf x}(t))\leq\mathcal{L}({\bf x}(t_{0}))\exp\left(-\min\left(c_{r},c_{\bf u}c_{\bf x}^{2}|I|\right)2Lc_{\bf x}^{2}\sigma_{\min}^{2}({\bf A}|_{I})(t-t_{0})\right)$ (3.4)
@@
See Appendix A. ∎
@@ 3.2 Invariants of the flow
Let $L\in\mathbb{N}$. Suppose ${\bf x}$ follows the dynamics in (1.11). Then the quantity
$h_{0}(t):=\left(I-{\bf A}^{\dagger}{\bf A}\right)\cdot\begin{cases}\log({\bf x}(t))&\text{if }L=2\\ {\bf x}(t)^{\odot 2-L}&\text{if }L\neq 2\end{cases}$ (3.5)
@@
Let $L\in\mathbb{N}$, ${\tilde{\eta}}>0$, and $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$. Suppose $r,{\bf u}$ follow the dynamics in (1.13) and (1.14) with $\|{\bf u}_{0}\|_{2}=1$. Then the quantity
$h_{\tilde{\eta}}(t):=(I-{\bf A}^{\dagger}{\bf A})\cdot\begin{cases}\log\left({\bf u}(t)\exp\left(\frac{1}{2{\tilde{\eta}}}r(t)^{2}\right)\right)&\text{if }L=2\\ {\bf u}(t)^{\odot 2-L}\exp\left(\frac{2-L}{2{\tilde{\eta}}}r(t)^{2}\right)&\text{if }L\neq 2\end{cases}$ (3.6)
@@
Proof of Lemma 3.8.
Since $\|{\bf u}_{0}\|_{2}=1$, by Lemma 3.2 $\|{\bf u}(t)\|_{2}=1$ for all $t\geq 0$. By substituting $\eta_{r}={\tilde{\eta}}$, $\eta_{\bf u}=1$, and $\|{\bf u}\|_{2}=1$ into the dynamics (A.5), we obtain
$\partial_{t}{\bf u}=-r\nabla\mathcal{L}\left(r{\bf u}\right)-\frac{r}{{\tilde{\eta}}}{\bf u}\partial_{t}r=-r\nabla\mathcal{L}\left(r{\bf u}\right)-\frac{1}{2{\tilde{\eta}}}\partial_{t}(r^{2}){\bf u}.$ (3.7)
@@
$(I-\mathcal{P}_{{\bf A}})\left({\bf u}^{\odot 1-L}\odot\nabla\mathcal{L}\left(r{\bf u}\right)\right)=(I-\mathcal{P}_{{\bf A}})\left({\bf A}^{\top}\left({\bf A}\left(r{\bf u}\right)^{\odot L}-{\bf b}\right)r^{L-1}\right)=0,$ (3.8)
by applying the operation $(I-\mathcal{P}_{{\bf A}})[{\bf u}^{\odot 1-L}\odot\,\cdot\,]$ to (3.7) we have
$(I-\mathcal{P}_{{\bf A}})({\bf u}^{\odot 1-L}\odot\partial_{t}{\bf u})=-\frac{1}{2{\tilde{\eta}}}\partial_{t}(r^{2})(I-\mathcal{P}_{{\bf A}}){\bf u}^{\odot 2-L}.$ (3.9)
@@
For $L=2$, since ${\bf u}^{\odot-1}\odot\partial_{t}{\bf u}=\partial_{t}\log({\bf u})$, (3.9) can be expressed as
@@
$(I-\mathcal{P}_{{\bf A}})\left(\log({\bf u})+\frac{r^{2}}{2{\tilde{\eta}}}\bm{1}\right)=(I-\mathcal{P}_{{\bf A}})\left(\log({\bf u}_{0})+\frac{r_{0}^{2}}{2{\tilde{\eta}}}\bm{1}\right).$ (3.10)
@@
For $L\neq 2$ the left hand side of (3.9) can be written as $\frac{1}{2-L}\partial_{t}(I-\mathcal{P}_{{\bf A}}){\bf u}^{\odot 2-L}$. Let $\widetilde{{\bf u}}=(I-\mathcal{P}_{{\bf A}}){\bf u}^{\odot 2-L}$. Then we have that
@@
$\widetilde{{\bf u}}\cdot\exp\left(\frac{2-L}{2{\tilde{\eta}}}r^{2}\right)=\widetilde{{\bf u}}_{0}\cdot\exp\left(\frac{2-L}{2{\tilde{\eta}}}r_{0}^{2}\right)$ (3.11)
@@
Let $L\in\mathbb{N}$ and ${\tilde{\eta}}>0$. Suppose ${\bf x}$ follows the dynamics in (1.11), and $(r,{\bf u})$ follow the dynamics in (1.13) and (1.14) with $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$, $r_{0},{\bf u}_{0}>0$, and $\|{\bf u}_{0}\|_{2}=1$. Denote ${\bf x}_{\text{wn}}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Suppose $r_{\infty}=\lim_{t\to\infty}r(t)$ exists and not equal to zero. Then for $L=2$,
$\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)\log({\bf x}(t))$ $=\left(I-\mathcal{P}_{{\bf A}}\right)\log({\bf x}(0))$ (3.12)
$\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right)\log({\bf x}_{\text{wn}}(t))$ $=(I-\mathcal{P}_{{\bf A}})\log\left(\gamma(r_{0},r_{\infty})\cdot{\bf x}_{\text{wn}}(0)\right)$ (3.13)
@@
$\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right){\bf x}^{\odot 2-L}(t)$ $=\left(I-\mathcal{P}_{{\bf A}}\right){\bf x}^{\odot 2-L}(0)$ (3.14)
$\lim_{t\to\infty}\left(I-\mathcal{P}_{{\bf A}}\right){\bf x}_{\text{wn}}^{\odot 2-L}(t)$ $=\left(I-\mathcal{P}_{{\bf A}}\right)[\,\gamma(r_{0},r_{\infty})\cdot{\bf x}_{\text{wn}}^{\odot L}(0)]^{\odot 2-L}$ (3.15)
@@
$\gamma(r_{0},r)$ $:=\frac{r}{r_{0}}\exp\left(\frac{r_{0}^{2}-r^{2}}{2{\tilde{\eta}}}\right).$ (3.16)
@@
Consider the same setting as in Lemma 3.9. Suppose that there exists ${\bf v}>0$ such ${\bf A}{\bf v}=0$ and that ${\bf x}_{\text{wn}},{\bf x}$ are bounded above. Then each entry of ${\bf x}_{\text{wn}}$ and ${\bf x}$ is also bounded away from zero.